BMC is considering upgrading the sound systems in their theaters so that they can patrons can get the full experience from surround sound movies. They discovered that upgrade costs at loca±ons with 12 screens were $170,000 but were $110,000 at loca±ons with six screens. The Fxed costs of upgrading at a loca±on are
1
Expert's answer
2017-02-01T13:33:14-0500

If there was discovered that upgrade costs at locations with 12 screens were $170,000 but were $110,000 at locations with six screens, then the fixed costs of upgrading at a location are: FC = 170,000 - 12x = 110,000 - 6x,
170,000 - 12x = 110,000 - 6x,
6x = 60,000,
x = $10,000 - the variable cost of one unit produced.
FC = 110,000 - 6*10,000 = $50,000.